OpenStudy (itsmichelle29):

Which of the following statements have the same result? Explain each step in solving each one. I.f(2) when f(x) = 3x + 2 II.f-1(3) when f(x) = 2 x minus 7, all over 3 III.2y + 14 = 4y - 2

OpenStudy (freckles):

did you replace x with 2 for I?

OpenStudy (freckles):

in II how you have to do is replace f(x) with 3 and solve for x. III. you solve for y.

OpenStudy (itsmichelle29):

so I is 8

OpenStudy (itsmichelle29):

@freckles

OpenStudy (itsmichelle29):

and II. i get -1/3 is that correct

OpenStudy (freckles):

let see 3(2)+2 is 6+2 which is 8 so yes for I

OpenStudy (freckles):

\[3=\frac{2x-7}{3}\] I think this is what you have above?

OpenStudy (freckles):

and I asked you to replace f(x) with 3 since we want to find f^(-1)(3)

OpenStudy (freckles):

hmm solving that shouldn't give -1/3

OpenStudy (itsmichelle29):

idk what the answer is

OpenStudy (itsmichelle29):

@freckles

OpenStudy (freckles):

did you try solving the equation I have above?

OpenStudy (freckles):

what is 3(3)?

OpenStudy (freckles):

\[3=\frac{2x-7}{3} \\ 3(3)=2x-7\]

OpenStudy (freckles):

and I will give you another hint add 7 on both sides

OpenStudy (itsmichelle29):

so its 8

OpenStudy (itsmichelle29):

so they all have the same result

OpenStudy (freckles):

it seems so :)